OpenStudy (anonymous):

i need to find the integral of (3x+2)^2 using substitution

OpenStudy (mimi_x3):

\[\int (3x+2)^{2}dx\] u = 3x + 2 => du / dx = 3 => \[\frac{du}{dx} = 3=> du = 3dx => 3dx = du => dx = du/3 \] \[\int (u)^2* \frac{du}{3} =>\int \frac{1}{3} u^{2}du\]

OpenStudy (mimi_x3):

\[ => \frac{1}{3} \int u^{2}du\]

OpenStudy (anonymous):

thanks

OpenStudy (mimi_x3):

:)